# FoundationPose: Unified 6D Pose Estimation and Tracking of Novel Objects

### v1

@@ Abstract
We present FoundationPose, a unified foundation model for 6D object pose estimation and tracking, supporting both model-based and model-free setups. Our approach can be instantly applied at test-time to a novel object without fine-tuning, as long as its CAD model is given, or a small number of reference images are captured. We bridge the gap between these two setups with a neural implicit representation that allows for effective novel view synthesis, keeping the downstream pose estimation modules invariant under the same unified framework. Strong generalizability is achieved via large-scale synthetic training, aided by a large language model (LLM), a novel transformer-based architecture, and contrastive learning formulation. Extensive evaluation on multiple public datasets involving challenging scenarios and objects indicate our unified approach outperforms existing methods specialized for each task by a large margin. In addition, it even achieves comparable results to instance-level methods despite the reduced assumptions. Project page: https://nvlabs.github.io/FoundationPose/
@@ 1 Introduction
Computing the rigid 6D transformation from the object to the camera, also known as object pose estimation, is crucial for a variety of applications, such as robotic manipulation [28, 60] and mixed reality [39]. Classic methods [19, 18, 45, 29] are known as instance-level because they only work on the specific object instance determined at training time. Such methods usually require a textured CAD model for generating training data, and they cannot be applied to an unseen novel object at test time. While category-level methods [56, 52, 5, 65, 31] remove these assumptions (instance-wise training and CAD models), they are limited to objects within the predefined category on which they are trained. Moreover, obtaining category-level training data is notoriously difficult, in part due to additional pose canonicalization and examination steps [56] that must be applied.
@@
In this paper we propose a unified framework called FoundationPose that performs both pose estimation and tracking for novel objects in both the model-based and model-free setups. As seen in Fig. 1, our method outperforms existing state-of-art methods specialized for each of these four tasks. Our strong generalizability is achieved via large-scale synthetic training, aided by a large language model (LLM), as well as a novel transformer-based architecture and contrastive learning. We bridge the gap between model-based and model-free setups with a neural implicit representation that allows for effective novel view synthesis with a small number (16) of reference images, achieving rendering speeds that are significantly faster than previous render-and-compare methods [33, 30, 59]. Our contributions can be summarized as follows:
@@ 2 Related Work
Object Pose Tracking. 6D object pose tracking aims to leverage temporal cues to enable more efficient, smooth and accurate pose prediction on video sequence. Through neural rendering, our method can be trivially extended to the pose tracking task with high efficiency. Similar to single-frame pose estimation, existing tracking methods can be categorized into their counterparts depending on the assumptions of object knowledge. These include instance-level methods [59, 33, 8], category-level methods [55, 36], model-based novel object tracking [49, 62, 27] and model-free novel object tracking [58, 61]. Under both model-based and model-free setups, we set a new benchmark record across public datasets, even outperforming state-of-art methods that require instance-level training [59, 33, 8].
@@ 3 Approach
Our method is described in the following subsections. The relationships between the subsections, and the system as a whole, are illustrated in Fig. 2.
@@ 3.1 Language-aided Data Generation at Scale
3.2 Neural Unknown Object Modeling
In the absence of CAD models, one key challenge is to represent the object to effectively render images with sufficient quality for downstream modules. Neural implicit representations have not only been shown to be effective for novel view synthesis, but they can also be parallelized on a GPU, thus providing high computational efficiency when rendering multiple pose hypotheses for downstream pose estimation modules. To this end, we introduce an object-centric neural SDF representation for object modeling, inspired by previous work [57, 41, 64, 61].
Field Representation. We represent the object by two functions [64] as shown in Fig. 2. First, the geometry function takes as input a 3D point and outputs a signed distance value . Second, the appearance function takes the intermediate feature vector from the geometry network, a point normal , and a view direction , and outputs the color . In practice, we apply multi-resolution hash encoding [41] to before forwarding to the network. Both and are embedded by a fixed set of second-order spherical harmonic coefficients. The implicit object surface is obtained by taking the zero level set of the signed distance field: . Compared to NeRF [40], the SDF representation provides higher quality depth rendering while removing the need to manually select a density threshold.
@@ 3.2 Neural Unknown Object Modeling
The learning is optimized per object without priors and can be efficiently performed within seconds. When training the pose refinement (Sec. 3.3) and selection (Sec. 3.4) modules, we first pretrain the neural object field with randomized number of synthetic reference images capturing the 3D asset. The trained neural object field then provides rendering which will be mixed with the model-based OpenGL rendering as input to subsequent networks. This better covers the distribution of both model-based and model-free setups, enabling effective generalization as a unified framework.
Rendering. The Neural Field only needs to be trained once for a novel unknown object. Once trained, the field is efficiently rendered at inference, serving as a drop-in replacement for a conventional graphics pipeline. In addition to the color rendering as in the original NeRF [40], we also need depth rendering for our RGBD based pose estimation and tracking. To do so, we perform marching cubes [38] to extract a mesh from the zero level set of the SDF. This only needs to be performed once for each object. At inference, given an object pose, we then render the depth image following the rasterization process. Alternatively, one could directly render the depth image using online with sphere tracing [13]; however, we found this leads to less efficiency, especially when there is a large number of pose hypotheses to render in parallel.
@@ 3.3 Pose Hypothesis Generation
Pose Initialization. Given the RGBD image, we first initialize the translation using the 3D point located at the median depth within the region of interest defined by the 2D detection. To initialize rotations, we uniformly sample viewpoints from an icosphere centered on the object with the camera facing the center. These camera poses are further augmented with discretized in-plane rotations, resulting in global pose initializations which are sent as input to the pose refiner.
@@
The refinement network architecture is illustrated in Fig. 2; details are in the appendix. We first extract feature maps from the two RGBD input branches with a single shared CNN encoder. The feature maps are concatenated, fed into CNN blocks with residual connection [15], and tokenized by dividing into patches [9] with position embedding. Finally, the network predicts the translation update and rotation update , each individually processed by a transformer encoder [54] and linearly projected to the output dimension. More concretely, represents the object’s translation shift in the camera frame, represents the object’s orientation update expressed in the camera frame. In practice, the rotations are parameterized with the 6D representation [68]. The input coarse pose is then updated by:
@@
where and are ground truth.
@@ 3.4 Pose Selection
Given a list of refined pose hypotheses, we build a pose ranking network to compute their scores. The pose with the highest score is selected as the final estimate.
Hierarchical Comparison. We propose a two-level comparison strategy. First, for each pose hypothesis, we compare the rendered image against the cropped input observation, where the pose-conditioned cropping operation was introduced in Sec. 3.3. This comparison is performed with a pose ranking encoder, where we utilize the same backbone architecture for feature extraction as in the refinement network. The extracted features are concatenated, tokenized and forwarded to the multi-head self-attention module so as to better leverage the global context for comparison. The pose ranking encoder performs average pooling to output a feature embedding describing the alignment quality between the rendering and the observation (Fig. 2 bottom-middle). At this point, we could directly project to a similarity scalar as typically done [42, 30, 2]. However, this would ignore the other pose hypotheses, forcing the network to output an absolute score assignment which can be difficult to learn.
Intuitively, we would like the network to leverage the global context of all pose hypotheses in order to make a more informed decision. Therefore, we introduce the second level of comparison among all the pose hypotheses, or more precisely, the concatenated feature embedding , which has encoded the pose alignment information. To adapt to varying , we treat as a sequence and perform multi-head self-attention which naturally generalizes to varying lengths [54]. We also do not apply position encoding to here, so as to be agnostic to the permutation. The attended feature is then linearly projected to the scores to be assigned to each pose hypothesis. The effectivenss of this hierarchical comparison strategy is shown in a typical example in Fig. 4.
@@ 4.1 Dataset and Setup
We consider 5 datasets: LINEMOD [21], Occluded LINEMOD [1], YCB-Video [63], T-LESS [23], and YCBInEOAT [59]. These involve various challenging scenarios (dense clutter, multi-instance, static or dynamic scenes, table-top or robotic manipulation), and objects with diverse properties (textureless, shiny, symmetric, varying sizes).
@@ 4.3 Pose Estimation Comparison
| Method |
Unseen |
Dataset | Mean | ||
@@
Model-based. Table 3 presents the comparison results among RGBD methods on 3 core datasets from BOP: Occluded-LINEMOD [1], YCB-Video [63] and T-LESS [23]. All methods use Mask R-CNN [16] for 2D detection. Our method outperforms the existing model-based methods that deal with novel objects by a large margin, including the instance-level method [14].
@@ 4.4 Pose Tracking Comparison
Solely evaluating on table-top static scenes does not expose challenges of abrupt out-of-plane rotations, dynamic external occlusions and disentangled camera motions [59]. Thus, for more comprehensive comparison, we also evaluate pose tracking methods on the YCBInEOAT [59] dataset which includes videos of dynamic robotic manipulation. Results under the model-based setup are presented in Table 4. Our method achieves the best performance and even outperforms the instance-wise training method [59] with ground-truth pose initialization. Moreover, our unified framework also allows for end-to-end pose estimation and tracking without external pose initialization, which is the only method with such capability, noted as Ours† in the table.
@@
| Re-initialization | Yes (290) | No | Yes (2) | No | No | No | No | No | |||||||||
| Novel object | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||||||
| Object setup | Model-based | Model-based | Model-based | Model-based | Model-based | Model-based | Model-based | Model-free | |||||||||
@@
| 003_cracker_box | 88.5 | 93.0 | 96.5 | 97.2 | 96.0 | 97.1 | 96.4 | 97.2 | 57.0 | 72.3 | 82.4 | 92.1 | 96.9 | 97.8 | 96.2 | 97.5 | |
| 004_sugar_box | 94.3 | 96.3 | 97.6 | 98.1 | 94.0 | 96.4 | 97.1 | 97.9 | 50.4 | 72.7 | 96.1 | 98.4 | 96.9 | 98.2 | 94.5 | 97.4 | |
| 005_tomato_soup_can | 89.1 | 93.2 | 95.0 | 97.2 | 87.2 | 95.2 | 64.7 | 89.5 | 72.4 | 91.6 | 73.2 | 97.3 | 96.3 | 98.1 | 94.3 | 97.9 | |
| 006_mustard_bottle | 92.0 | 95.1 | 95.8 | 97.4 | 98.3 | 98.5 | 97.1 | 98.0 | 87.7 | 98.2 | 96.2 | 98.4 | 97.3 | 98.4 | 97.3 | 98.5 | |
| 007_tuna_fish_can | 92.0 | 96.4 | 86.5 | 91.1 | 86.8 | 93.6 | 69.1 | 93.3 | 28.7 | 52.9 | 73.2 | 95.8 | 96.9 | 98.5 | 84.0 | 97.8 | |
| 008_pudding_box | 80.1 | 88.3 | 97.9 | 98.4 | 60.9 | 87.1 | 96.8 | 97.9 | 12.7 | 18.0 | 73.8 | 88.9 | 97.8 | 98.5 | 96.9 | 98.5 | |
| 009_gelatin_box | 92.0 | 94.4 | 97.8 | 98.4 | 98.2 | 98.6 | 97.5 | 98.4 | 49.1 | 70.7 | 97.2 | 98.8 | 97.7 | 98.5 | 97.6 | 98.5 | |
| 010_potted_meat_can | 78.0 | 88.9 | 77.8 | 84.2 | 76.4 | 83.5 | 83.7 | 86.7 | 44.1 | 45.6 | 93.3 | 97.3 | 95.1 | 97.7 | 94.8 | 97.5 | |
| 011_banana | 81.0 | 90.5 | 94.9 | 97.2 | 92.8 | 97.7 | 86.3 | 96.1 | 93.3 | 97.7 | 95.6 | 98.4 | 96.4 | 98.4 | 95.6 | 98.1 | |
| 019_pitcher_base | 90.4 | 94.7 | 96.8 | 97.5 | 97.7 | 98.1 | 97.3 | 97.7 | 97.9 | 98.2 | 97.0 | 98.8 | 96.7 | 98.0 | 96.8 | 98.0 | |
| 021_bleach_cleanser | 81.7 | 90.5 | 95.9 | 97.2 | 95.9 | 97.0 | 95.2 | 97.2 | 95.9 | 97.3 | 92.6 | 97.5 | 95.5 | 97.8 | 94.7 | 97.5 | |
| 024_bowl | 38.8 | 90.6 | 80.9 | 94.5 | 34.0 | 93.0 | 30.4 | 97.2 | 24.2 | 82.4 | 74.4 | 98.4 | 95.2 | 97.6 | 90.5 | 95.3 | |
| 025_mug | 83.2 | 92.0 | 91.5 | 96.9 | 86.9 | 96.7 | 83.2 | 93.3 | 60.0 | 71.2 | 95.6 | 98.5 | 95.6 | 97.9 | 91.5 | 96.1 | |
| 035_power_drill | 85.4 | 92.3 | 96.4 | 97.4 | 97.8 | 98.2 | 97.1 | 97.8 | 97.9 | 98.3 | 96.7 | 98.5 | 96.9 | 98.2 | 96.3 | 97.9 | |
| 036_wood_block | 44.3 | 75.4 | 95.2 | 96.7 | 37.8 | 93.6 | 95.5 | 96.9 | 45.7 | 62.5 | 93.5 | 97.2 | 93.2 | 97.0 | 92.9 | 97.0 | |
| 037_scissors | 70.3 | 84.5 | 95.7 | 97s | 72.7 | 85.5 | 4.2 | 16.2 | 20.9 | 38.6 | 93.5 | 97.3 | 94.8 | 97.5 | 95.5 | 97.8 | |
| 040_large_marker | 80.4 | 91.2 | 92.2 | 96.0 | 89.2 | 97.3 | 35.6 | 53.0 | 12.2 | 18.9 | 88.5 | 97.8 | 96.9 | 98.6 | 96.6 | 98.6 | |
| 051_large_clamp | 73.9 | 84.1 | 94.7 | 96.9 | 90.1 | 95.5 | 61.2 | 72.3 | 62.8 | 80.1 | 91.8 | 96.9 | 93.6 | 97.3 | 92.5 | 96.7 | |
| 052_extra_large_clamp | 49.3 | 90.3 | 91.7 | 95.8 | 84.4 | 94.1 | 93.7 | 96.6 | 67.5 | 69.7 | 85.9 | 94.3 | 94.4 | 97.5 | 93.4 | 97.3 | |

### v2

@@ Abstract
We present FoundationPose, a unified foundation model for 6D object pose estimation and tracking, supporting both model-based and model-free setups. Our approach can be instantly applied at test-time to a novel object without fine-tuning, as long as its CAD model is given, or a small number of reference images are captured. Thanks to the unified framework, the downstream pose estimation modules are the same in both setups, with a neural implicit representation used for efficient novel view synthesis when no CAD model is available. Strong generalizability is achieved via large-scale synthetic training, aided by a large language model (LLM), a novel transformer-based architecture, and contrastive learning formulation. Extensive evaluation on multiple public datasets involving challenging scenarios and objects indicate our unified approach outperforms existing methods specialized for each task by a large margin. In addition, it even achieves comparable results to instance-level methods despite the reduced assumptions. Project page: https://nvlabs.github.io/FoundationPose/
@@ 1 Introduction
Computing the rigid 6D transformation from the object to the camera, also known as object pose estimation, is crucial for a variety of applications, such as robotic manipulation [30, 69, 70] and mixed reality [43]. Classic methods [21, 20, 50, 68, 31] are known as instance-level because they only work on the specific object instance determined at training time. Such methods usually require a textured CAD model for generating training data, and they cannot be applied to an unseen novel object at test time. While category-level methods [64, 60, 5, 75, 34] remove these assumptions (instance-wise training and CAD models), they are limited to objects within the predefined category on which they are trained. Moreover, obtaining category-level training data is notoriously difficult, in part due to additional pose canonicalization and examination steps [64] that must be applied.
@@
In this paper we propose a unified framework called FoundationPose that performs both pose estimation and tracking for novel objects in both the model-based and model-free setups, using RGBD images. As seen in Fig. 1, our method outperforms existing state-of-art methods specialized for each of these four tasks. Our strong generalizability is achieved via large-scale synthetic training, aided by a large language model (LLM), as well as a novel transformer-based architecture and contrastive learning. We bridge the gap between model-based and model-free setups with a neural implicit representation that allows for effective novel view synthesis with a small number (16) of reference images, achieving rendering speeds that are significantly faster than previous render-and-compare methods [36, 32, 67]. Our contributions can be summarized as follows:
@@ 2 Related Work
Object Pose Tracking. 6D object pose tracking aims to leverage temporal cues to enable more efficient, smooth and accurate pose prediction on video sequence. Through neural rendering, our method can be trivially extended to the pose tracking task with high efficiency. Similar to single-frame pose estimation, existing tracking methods can be categorized into their counterparts depending on the assumptions of object knowledge. These include instance-level methods [67, 36, 8, 11], category-level methods [63, 39], model-based novel object tracking [56, 72, 29] and model-free novel object tracking [66, 71]. Under both model-based and model-free setups, we set a new benchmark record across public datasets, even outperforming state-of-art methods that require instance-level training [67, 36, 8].
@@ 3 Approach
Our system as a whole is illustrated in Fig. 2, showing the relationships between the various components, which are described in the following subsections.
@@ 3.1 Language-aided Data Generation at Scale
3.2 Neural Object Modeling
For the model-free setup, when the 3D CAD model is unavailable, one key challenge is to represent the object to effectively render images with sufficient quality for downstream modules. Neural implicit representations are both effective for novel view synthesis and parallelizable on a GPU, thus providing high computational efficiency when rendering multiple pose hypotheses for downstream pose estimation modules, as shown in Fig. 2. To this end, we introduce an object-centric neural field representation for object modeling, inspired by previous work [65, 45, 74, 71].
Field Representation. We represent the object by two functions [74] as shown in Fig. 2. First, the geometry function takes as input a 3D point and outputs a signed distance value . Second, the appearance function takes the intermediate feature vector from the geometry network, a point normal , and a view direction , and outputs the color . In practice, we apply multi-resolution hash encoding [45] to before forwarding to the network. Both and are embedded by a fixed set of second-order spherical harmonic coefficients. The implicit object surface is obtained by taking the zero level set of the signed distance field (SDF): . Compared to NeRF [44], the SDF representation provides higher quality depth rendering while removing the need to manually select a density threshold.
@@ 3.2 Neural Object Modeling
The learning is optimized per object without priors and can be efficiently performed within seconds. The neural field only needs to be trained once for a novel object.
Rendering. Once trained, the neural field can be used as a drop-in replacement for a conventional graphics pipeline, to perform efficient rendering of the object for subsequent render-and-compare iterations. In addition to the color rendering as in the original NeRF [44], we also need depth rendering for our RGBD based pose estimation and tracking. To do so, we perform marching cubes [41] to extract a textured mesh from the zero level set of the SDF, combined with color projection. This only needs to be performed once for each object. At inference, given an object pose, we then render the RGBD image following the rasterization process. Alternatively, one could directly render the depth image using online with sphere tracing [14]; however, we found this leads to less efficiency, especially when there is a large number of pose hypotheses to render in parallel.
@@ 3.3 Pose Hypothesis Generation
Pose Initialization. Given the RGBD image, the object is detected using an off-the-shelf method such as Mask R-CNN [18] or CNOS [47]. We initialize the translation using the 3D point located at the median depth within the detected 2D bounding box. To initialize rotations, we uniformly sample viewpoints from an icosphere centered on the object with the camera facing the center. These camera poses are further augmented with discretized in-plane rotations, resulting in global pose initializations which are sent as input to the pose refiner.
@@
The refinement network architecture is illustrated in Fig. 2; details are in the appendix. We first extract feature maps from the two RGBD input branches with a single shared CNN encoder. The feature maps are concatenated, fed into CNN blocks with residual connection [17], and tokenized by dividing into patches [9] with position embedding. Finally, the network predicts the translation update and rotation update , each individually processed by a transformer encoder [62] and linearly projected to the output dimension. More concretely, represents the object’s translation shift in the camera frame, represents the object’s orientation update expressed in the camera frame. In practice, the rotations are parameterized in axis-angle representation. We also experimented with the 6D representation [78] which achieves similar results. The input coarse pose is then updated by:
@@
where and are ground truth; and are the weights balancing the losses, which are set to 1 empirically.
@@ 3.4 Pose Selection
Given a list of refined pose hypotheses, we use a hierarchical pose ranking network to compute their scores. The pose with the highest score is selected as the final estimate.
Hierarchical Comparison. The network uses a two-level comparison strategy. First, for each pose hypothesis, the rendered image is compared against the cropped input observation, using the pose-conditioned cropping operation was introduced in Sec. 3.3. This comparison (Fig. 2 bottom-left) is performed with a pose ranking encoder, utilizing the same backbone architecture for feature extraction as in the refinement network. The extracted features are concatenated, tokenized and forwarded to the multi-head self-attention module so as to better leverage the global image context for comparison. The pose ranking encoder performs average pooling to output a feature embedding describing the alignment quality between the rendering and the observation (Fig. 2 bottom-middle). At this point, we could directly project to a similarity scalar as typically done [46, 32, 2]. However, this would ignore the other pose hypotheses, forcing the network to output an absolute score assignment which can be difficult to learn.
To leverage the global context of all pose hypotheses in order to make a more informed decision, we introduce a second level of comparison among all the pose hypotheses. Multi-head self-attention is performed on the concatenated feature embedding , which encodes the pose alignment information from all poses. By treating as a sequence, this approach naturally generalizes to varying lengths of [62]. We do not apply position encoding to , so as to be agnostic to the permutation. The attended feature is then linearly projected to the scores to be assigned to the pose hypotheses. The effectiveness of this hierarchical comparison strategy is shown in a typical example in Fig. 4.
@@ 4.1 Dataset and Setup
We consider 5 datasets: LINEMOD [23], Occluded-LINEMOD [1], YCB-Video [73], T-LESS [25], and YCBInEOAT [67]. These involve various challenging scenarios (dense clutter, multi-instance, static or dynamic scenes, table-top or robotic manipulation), and objects with diverse properties (textureless, shiny, symmetric, varying sizes).
@@ 4.3 Pose Estimation Comparison
Method | Unseen | Dataset | Mean | ||
@@
Model-based. Table 3 presents the comparison results among RGBD methods on 3 core datasets from BOP: Occluded-LINEMOD [1], YCB-Video [73] and T-LESS [25]. All methods use Mask R-CNN [18] for 2D detection. Our method outperforms the existing model-based methods that deal with novel objects, and the instance-level method [15], by a large margin.
@@ 4.4 Pose Tracking Comparison
For comprehensive comparison on the challenges of abrupt out-of-plane rotations, dynamic external occlusions and disentangled camera motions, we evaluate pose tracking methods on the YCBInEOAT [67] dataset which includes videos of dynamic robotic manipulation. Results under the model-based setup are presented in Table 4. Our method achieves the best performance and even outperforms the instance-wise training method [67] with ground-truth pose initialization. Moreover, our unified framework also allows for end-to-end pose estimation and tracking without external pose initialization, which is the only method with such capability, noted as Ours† in the table.
@@
Re-initialization | Yes (290) | No | Yes (2) | No | No | No | No | No | ||||||||||||
Novel object | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
Object setup | Model-based | Model-based | Model-based | Model-based | Model-based | Model-based | Model-based | Model-free | ||||||||||||
@@
003_cracker_box | 88.5 | 93.0 | 96.5 | 97.2 | 96.0 | 97.1 | 96.4 | 97.2 | 57.0 | 72.3 | 82.4 | 92.1 | 96.9 | 97.8 | 96.2 | 97.5 | ||||
004_sugar_box | 94.3 | 96.3 | 97.6 | 98.1 | 94.0 | 96.4 | 97.1 | 97.9 | 50.4 | 72.7 | 96.1 | 98.4 | 96.9 | 98.2 | 94.5 | 97.4 | ||||
005_tomato_soup_can | 89.1 | 93.2 | 95.0 | 97.2 | 87.2 | 95.2 | 64.7 | 89.5 | 72.4 | 91.6 | 73.2 | 97.3 | 96.3 | 98.1 | 94.3 | 97.9 | ||||
006_mustard_bottle | 92.0 | 95.1 | 95.8 | 97.4 | 98.3 | 98.5 | 97.1 | 98.0 | 87.7 | 98.2 | 96.2 | 98.4 | 97.3 | 98.4 | 97.3 | 98.5 | ||||
007_tuna_fish_can | 92.0 | 96.4 | 86.5 | 91.1 | 86.8 | 93.6 | 69.1 | 93.3 | 28.7 | 52.9 | 73.2 | 95.8 | 96.9 | 98.5 | 84.0 | 97.8 | ||||
008_pudding_box | 80.1 | 88.3 | 97.9 | 98.4 | 60.9 | 87.1 | 96.8 | 97.9 | 12.7 | 18.0 | 73.8 | 88.9 | 97.8 | 98.5 | 96.9 | 98.5 | ||||
009_gelatin_box | 92.0 | 94.4 | 97.8 | 98.4 | 98.2 | 98.6 | 97.5 | 98.4 | 49.1 | 70.7 | 97.2 | 98.8 | 97.7 | 98.5 | 97.6 | 98.5 | ||||
010_potted_meat_can | 78.0 | 88.9 | 77.8 | 84.2 | 76.4 | 83.5 | 83.7 | 86.7 | 44.1 | 45.6 | 93.3 | 97.3 | 95.1 | 97.7 | 94.8 | 97.5 | ||||
011_banana | 81.0 | 90.5 | 94.9 | 97.2 | 92.8 | 97.7 | 86.3 | 96.1 | 93.3 | 97.7 | 95.6 | 98.4 | 96.4 | 98.4 | 95.6 | 98.1 | ||||
019_pitcher_base | 90.4 | 94.7 | 96.8 | 97.5 | 97.7 | 98.1 | 97.3 | 97.7 | 97.9 | 98.2 | 97.0 | 98.8 | 96.7 | 98.0 | 96.8 | 98.0 | ||||
021_bleach_cleanser | 81.7 | 90.5 | 95.9 | 97.2 | 95.9 | 97.0 | 95.2 | 97.2 | 95.9 | 97.3 | 92.6 | 97.5 | 95.5 | 97.8 | 94.7 | 97.5 | ||||
024_bowl | 38.8 | 90.6 | 80.9 | 94.5 | 34.0 | 93.0 | 30.4 | 97.2 | 24.2 | 82.4 | 74.4 | 98.4 | 95.2 | 97.6 | 90.5 | 95.3 | ||||
025_mug | 83.2 | 92.0 | 91.5 | 96.9 | 86.9 | 96.7 | 83.2 | 93.3 | 60.0 | 71.2 | 95.6 | 98.5 | 95.6 | 97.9 | 91.5 | 96.1 | ||||
035_power_drill | 85.4 | 92.3 | 96.4 | 97.4 | 97.8 | 98.2 | 97.1 | 97.8 | 97.9 | 98.3 | 96.7 | 98.5 | 96.9 | 98.2 | 96.3 | 97.9 | ||||
036_wood_block | 44.3 | 75.4 | 95.2 | 96.7 | 37.8 | 93.6 | 95.5 | 96.9 | 45.7 | 62.5 | 93.5 | 97.2 | 93.2 | 97.0 | 92.9 | 97.0 | ||||
037_scissors | 70.3 | 84.5 | 95.7 | 97s | 72.7 | 85.5 | 4.2 | 16.2 | 20.9 | 38.6 | 93.5 | 97.3 | 94.8 | 97.5 | 95.5 | 97.8 | ||||
040_large_marker | 80.4 | 91.2 | 92.2 | 96.0 | 89.2 | 97.3 | 35.6 | 53.0 | 12.2 | 18.9 | 88.5 | 97.8 | 96.9 | 98.6 | 96.6 | 98.6 | ||||
051_large_clamp | 73.9 | 84.1 | 94.7 | 96.9 | 90.1 | 95.5 | 61.2 | 72.3 | 62.8 | 80.1 | 91.8 | 96.9 | 93.6 | 97.3 | 92.5 | 96.7 | ||||
052_extra_large_clamp | 49.3 | 90.3 | 91.7 | 95.8 | 84.4 | 94.1 | 93.7 | 96.6 | 67.5 | 69.7 | 85.9 | 94.3 | 94.4 | 97.5 | 93.4 | 97.3 | ||||